### v1

@@ 2 Quantum mechanical closure framework
In this section we provide a brief overview of QMCL as it has been developed in [20, 16, 17]. The core of the QMCL framework consists of two steps: (1) embedding the classical dynamics into an infinite dimensional quantum mechanical system; (2) projecting the infinite dimensional system to a finite dimensional subsystem. The infinite dimensional embedding of the classical dynamics is treated in Section 2.1. In Section 2.2 we present the process used to project the embedded dynamics to a finite dimensional subsystem and derive a closure model.
In the description of QMCL provided below we intentionally omit some of the details required for its precise description and implementation. A detailed treatment of the scheme, including its data driven realization, will be given in the following sections once we extend it to the closure of dynamics governed by partial differential equations. A detailed analysis of QMCL as presented below can be found in [20, 16, 17].
@@ 2.1 Infinite dimensional embedding
To build the quantum mechanical closure framework we start by embedding the classical dynamics on state space $X$ into a quantum mechanical system on the infinite dimensional Hilbert space $H:=L^{2}(X,\mu;\mathbb{R})$. This involves embedding observables (functions of state) and probability densities, as well as the evolution of said densities under the dynamics and their bayesian conditioning.
@@
In the following we denote by $B(H)$ the space of bounded operators on Hilbert space $H$. We write $Q(H)$ to denote the space of quantum states on $H$, which are positive, trace class operators of unit trace. Note that positivity implies self adjointness in this context. Quantum states are also referred to as quantum density operators. In this work we mainly use pure quantum states, which are rank-1 quantum density operators and are equivalent to projections along a unit vector [23, 34, 36].
@@
Every classical observable $f\in L^{\infty}(X,\mu;\mathbb{R})\subset H$ is mapped to a self adjoint multiplication operator $A_{f}\in B(H)$, $A_{f}g=fg$ for all $g\in H$, referred to as a quantum observable. Additionally, every classical probability density $p\in L^{1}(X,\mu;\mathbb{R})$, $\sqrt{p}\in H$, is mapped to the quantum density operator $\rho_{p}\in Q(H)$, $\rho_{p}=\langle\sqrt{p},\cdot\rangle\sqrt{p}$, representing a pure quantum state. The expected value of observable $A\in B(H)$ when the quantum system is in state $\rho\in Q(H)$ is given by $\mathbb{E}_{\rho}A=\operatorname{tr}(\rho A)\in\mathbb{R}$, which is connected to the theory of quantum measurement [23, 34, 36].
@@
We use the notion of quantum measurement to define the surrogate flux map $\tilde{Z}:\tilde{\mathcal{Y}}\rightarrow\mathcal{Z}$ needed for the evolution of the resolved dynamics. In particular using a quantum state $\rho\in Q(H)$ to model the state of the unresolved degrees of freedom, and a finite number of operators $A^{(j)}\in B(H)$, $j\in\{0,\ldots,d-1\}$, to model $d\in\mathbb{N}$ observables of the system’s unresolved state, we define each component of the surrogate flux map $\tilde{Z}:Q(H)\rightarrow\mathbb{R}^{d}$ as
$\tilde{Z}^{(j)}(\rho)=\mathbb{E}_{\rho}(A^{(j)})=\operatorname{tr}(\rho A^{(j)}).$
@@
In other words we use flux space $\mathcal{Z}=\mathbb{R}^{d}$ and the space of quantum states $\tilde{\mathcal{Y}}=Q(H)$ to model the unresolved degrees of freedom of the dynamics.
@@
In addition to evaluating the required flux terms we need to be able to evolve a quantum state under the system’s dynamics and condition it based on observed dynamical behavior. In the classical case, the square root of probability densities $\sqrt{p}\in H$ evolves according to the unitary group of transfer operators, $P^{n}:H\rightarrow H$, $P^{n}\sqrt{p}=\sqrt{p}\circ\Phi^{-n}$, $n\in\mathbb{N}$. Accordingly, the evolution of quantum states $\rho\in Q(H)$ is governed by the induced transfer operator $\mathcal{P}^{n}:Q(H)\rightarrow Q(H)$, $\mathcal{P}^{n}\rho=P^{n}\rho P^{-n}$. In analogy to bayesian conditioning of classical probability densities, we can condition a quantum state $\rho\in Q(H)$ using a quantum effect operator $e\in B(H)$, $0\leq e\leq\operatorname{Id}$,
(2) $\rho|_{e}=\frac{\sqrt{e}\rho\sqrt{e}}{\mathbb{E}_{\rho}e},$
@@
A quantum effect operator can be thought of as the quantum analogue of an indicator function used in classical probability to represent an event, namely a set in the event space. For instance, for a given set $B\subset X$ we can build the indicator function $1_{B}\in H$ taking the value one on $B$ and zero elsewhere. The quantum effect operator $e_{B}\in B(H)$ can then be defined as multiplication by $1_{B}$, namely $e_{B}f=1_{B}f$ for all $f\in H$.
@@ 2.2 Closure modeling
We build a finite dimensional subspace $H^{L}\subset H$ using the leading $L$ elements of an orthonormal basis $\{\phi_{\ell}\}_{\ell=0}^{L-1}\subset H$. To obtain such a basis we define a kernel integral operator $K:H\rightarrow H$
@@
using a symmetric kernel function $k\in L^{2}(X\times X,\mu\times\mu;\mathbb{R})$. Operator $K$ is compact and self adjoint, so its eigenfunctions form an orthonormal basis of $H$. We define $H^{L}$ as the linear span of its leading $L$ eigenfunctions. Having built $H^{L}$, we project operators acting in $H$ to operators acting in $H^{L}$. This is done for all operators needed for the quantum mechanical embedding of the original classical dynamics.
@@
To build our closure modeling framework we set $\tilde{\mathcal{Y}}=Q(H^{L})$; namely, we use quantum states to model the unresolved degrees of freedom of the surrogate dynamics $\tilde{\Phi}^{t}$ from (1). Having made that choice, we now describe the calculation of fluxes needed for evolving the resolved state under $\tilde{\Phi}_{\mathcal{X}}^{t}$, and the surrogate map $\tilde{\Phi}_{\mathcal{Y}}^{t}$ used to evolve the quantum state representing the unresolved state. To do so, we use a discrete time version of the above flow maps, denoting them by $\tilde{\Phi}_{\mathcal{X}}^{n\Delta t}$ and $\tilde{\Phi}_{\mathcal{Y}}^{n\Delta t}$ with $n\in\mathbb{N}$ and timestep $\Delta t>0$.
@@
Given the resolved system state $\hat{x}_{i}\in\mathcal{X}$ at timestep $t_{i}$, we want to compute state $\hat{x}_{i+1}=\tilde{\Phi}_{\mathcal{X}}(\hat{x}_{i},\tilde{z}_{i})$ using flux terms $\tilde{z}_{i}\in\mathcal{Z}$. We use $\mathcal{Z}=\mathbb{R}^{d}$ for some $d\in\mathbb{N}$; that is, $\tilde{z}$ has components $\tilde{z}_{i}^{(j)}\in\mathbb{R}$, $j\in\{0,\ldots,d-1\}$. To calculate the component fluxes we use a quantum state $\rho_{i}\in Q(H^{L})$ representing the unresolved state at time $t_{i}$, and $d$ quantum observables $A^{(j)}\in B(H^{L})$. Each observable $A^{(j)}$ is a projected multiplication operator determined by the corresponding components $Z^{(j)}$ of the true flux function $Z:\mathcal{Y}\rightarrow\mathcal{Z}$. Intuitively, each $A^{(j)}$ represents the possible range of values of the corresponding true flux term $z^{(j)}$. Each surrogate flux term $\tilde{z}_{i}^{(j)}$ is computed as the expectation value of observable $A^{(j)}$ with respect to quantum state $\rho_{i}$
$\tilde{z}_{i}^{(j)}=\tilde{Z}^{(j)}(\rho_{i})=\mathbb{E}_{\rho_{i}}(A^{(j)})=\operatorname{tr}{(\rho_{i}A^{(j)})}.$
The resolved state is then updated by $x_{i+1}=\tilde{\Phi}_{\mathcal{X}}^{\Delta t}(x_{i},\tilde{z}_{i})$.
@@
Next we define the surrogate unresolved dynamics map $\tilde{\Phi}_{\mathcal{Y}}^{\Delta t}$ used to update the quantum state, $\rho_{i+1}=\tilde{\Phi}_{\mathcal{Y}}^{\Delta t}(\hat{x_{i}},\rho_{i})$. The quantum state update consists of two steps, a prediction and a correction step. Given the current quantum state $\rho_{i}$, we first use the projected transfer operator $\mathcal{P}_{L}:Q(H^{L})\rightarrow Q(H^{L})$ to compute the prior state
$\rho_{i+1}^{\prime}=\frac{\mathcal{P}_{L}\rho_{i}}{\operatorname{tr}{(\mathcal{P}_{L}\rho_{i})}}.$
@@
Renormalizing the quantum state after acting with $\mathcal{P}_{L}$ is required because the projected operator $\mathcal{P}_{L}$ is generally not unitary. Mapping $\rho_{i}\mapsto\rho_{i+1}^{\prime}$ corresponds to the prediction step. The correction step involves using the updated resolved state $\hat{x}_{i+1}$ to correct our prediction $\rho_{i+1}^{\prime}$. To do so, we use state $\hat{x}_{i+1}$ to build the projected quantum effect operator $e_{i+1}\in B(H^{L})$. Using the quantum Bayes rule (2) we condition the prior state $\rho_{i+1}^{\prime}$ to compute the posterior state
$\rho_{i+1}=\rho_{i+1}^{\prime}|_{e_{i+1}}=\frac{\sqrt{e_{i+1}}\rho_{i+1}^{\prime}\sqrt{e_{i+1}}}{\operatorname{tr}{(\sqrt{e_{i+1}}\rho_{i+1}^{\prime}\sqrt{e_{i+1}}})}.$
@@
The above two steps make up the surrogate unresolved dynamics $\tilde{\Phi}_{\mathcal{Y}}^{\Delta t}$. Although not unitary, operator $\mathcal{P}_{L}$ is trace nonincreasing, which means that it can be interpreted as governing the evolution of the open quantum system $H^{L}\subset H$. Under this light, the reduction of $\operatorname{tr}(\rho)$ when acting with $\mathcal{P}_{L}$ can be viewed as decoherence or leakage of density due to interaction with the environment [23, 34, 36].
@@
In the diagram $\operatorname{Id}$ denotes the formal identity map.
@@ 3 Closure of partial differential equations
In Section 3.1 we define the appropriate function space $H$ and describe the process of embedding the classical spatiotemporal dynamics into a quantum mechanical system in $H$. In Section 3.2 we introduce the kernel we use to discretize (compress) the embedded dynamics to a finite dimensional subsystem. In Sections 3.3 and 3.4 we explore the symmetry invariance properties of the derived eigenfunction basis and their consequences. In Sections 3.5 and 3.6 we develop the data driven formulation of the extended QMCL scheme which can be realized numerically.
@@ 3.1 Closure modeling
As explained in detail in Section 1.1, we consider spatiotemporal dynamics $\Phi^{t}:X\rightarrow X$, $t\geq 0$, possessing ergodic invariant measure $\mu$ with compact support, and infinite dimensional state space $X\subset L^{2}(S,\nu;\mathbb{R})$ where $S$ denotes a spatial domain. We assume the state space decomposition $X=\mathcal{X}\times\mathcal{Y}$ where $\mathcal{X}$ is the finite dimensional resolved state space and $\mathcal{Y}$ the unresolved state space.
To provide a general definition of the resolved state space $\mathcal{X}\subset X$ we consider measure space $(S,\mathcal{A},\nu)$ where $\mathcal{A}$ denotes the Borel $\sigma$-algebra on $S$. We form the subalgebra $\mathcal{A}_{M}\subset\mathcal{A}$ generated by a finite collection of $M\in\mathbb{N}$ disjoint subsets $S_{m}\subset S$, $m\in\{0,\ldots,M-1\}$, such that $\cup_{m=0}^{M-1}S_{m}=S$. Next we define measure $\nu_{M}$ as the restriction of $\nu$ to $\sigma$-algebra $\mathcal{A}_{M}$. The resolved state space is defined as the $M$-dimensional subspace $\mathcal{X}=L^{2}(S,\nu_{M};\mathbb{R})\cap X$. As a Hilbert space $\mathcal{X}$ is isomorphic to $\mathbb{R}^{M}$ with a scaled euclidean inner product. Each subset $S_{m}\subset S$ used to generate subalgebra $\mathcal{A}_{M}$ can also be represented by its center point $s_{m}\in S_{m}$. We denote by $S_{M}=\{s_{m}\}_{m=0}^{M-1}\subset S$ the collection of those center points.
To embed the classical dynamics $\Phi^{t}$ we define the Hilbert space of vector valued observables $H=L^{2}(X,\mu;H_{S})$ with $H_{S}=L^{2}(S,\nu_{M};\mathbb{R})$. Introducing product state space $\Omega=X\times S$ and measure $\sigma=\mu\times\nu_{M}$ we define $H_{\Omega}=L^{2}(\Omega,\sigma;\mathbb{R})$ and note the following Hilbert space isomorphisms
@@
where $H_{X}=L^{2}(X,\mu;\mathbb{R})$. Namely, for every vector valued observable $\vec{f}\in H$ there is a corresponding real valued $f\in H_{\Omega}$ such that for almost every $x\in X$ and $s\in S$ $f(x,\cdot)=\vec{f}(x)\in H_{S}$ and $f(x,s)=\vec{f}(x)(s)\in\mathbb{R}$. The motivation for our choice of space $H$ is our interest in observables that can effectively capture spatiotemporal dynamics; that is, patterns of evolution that have both a temporal dependence $x\in X$ and a spatial dependence $s\in S$. Given an observable $\vec{f}\in H$ the map $t\mapsto\vec{f}(\Phi^{t}(x))$ represents a spatiotemporal pattern generated by the dynamics $\Phi^{t}$ [22].
[…]
The definitions of quantum observables, quantum effects and the induced transfer operator carry over from Section 2 after using the new definition of Hilbert space $H$. Quantum states are now modeled as an operator valued field over $S$, $\rho\in L^{2}(S,\nu_{M};B(H))$ with $\operatorname{ran}{\rho}\subseteq Q(H)\subset B(H)$. This allows us to compute fluxes that vary over the spatial domain $S$ for each quantum observable $A^{(j)}\in B(H)$, $j\in\{0,\ldots,d-1\}$. More specifically, the surrogate flux map is defined as $\tilde{Z}:L^{2}(S,\nu_{M};B(H))\rightarrow L^{2}(S,\nu_{M};\mathbb{R}^{d})$ with action for each component
$\tilde{Z}^{(j)}(\rho)(s)=\operatorname{tr}(\rho(s)A^{(j)})\in\mathbb{R}$
where $\rho(s)\in Q(H)\subset B(H)$ denotes the quantum state representing point $s\in S$.
In the above construction we have made the space choices $\tilde{\mathcal{Y}}=L^{2}(S,\nu_{M};B(H))$ and $\mathcal{Z}=L^{2}(S,\nu_{M};\mathbb{R}^{d})$. Namely, we use one quantum state to represent the unresolved degrees of freedom of the dynamics at each point in the spatial domain $S$, which in turn generate $d$ flux terms at each spatial point using the quantum observables $A^{(j)}$.

### v2

@@ 3 Quantum mechanical closure framework
In this section we provide a brief overview of QMCl as it has been developed in [56, 57, 55], which is concerned with the closure of dynamics governed by ordinary differential equations. The core of the QMCl framework consists of two steps: (1) embedding the classical dynamics into an infinite dimensional quantum mechanical system; (2) projecting the infinite dimensional system to a finite dimensional subsystem. The infinite dimensional embedding of the classical dynamics is treated in Section 3.1. In Section 3.2 we present the process used to project the embedded dynamics to a finite dimensional subsystem and derive a closure model.
In the description of QMCl provided below we omit some of the details required for its precise description and implementation. A detailed treatment of the scheme, including its data driven realization, will be given in the following sections once we extend it to the closure of dynamics governed by partial differential equations. A detailed analysis of QMCl as presented below can be found in [56, 57, 55].
@@ 3.1 Infinite dimensional embedding
To build the quantum mechanical closure framework we start by embedding the classical dynamics on state space $X$ into a quantum mechanical system on the infinite dimensional Hilbert space $H=L^{2}(X,\mu;\mathbb{C})$. This involves embedding observables (functions of state) and probability densities, as well as the evolution of said densities under the dynamics and their bayesian conditioning.
@@
In the following we denote by $B(H)$ the space of bounded linear operators on Hilbert space $H$. We write $Q(H)$ to denote the space of quantum states on $H$ which are positive, trace class operators of unit trace. We use the term positive operator to describe a selfadjoint operator with nonnegative spectrum. Quantum states are also referred to as quantum density operators, and can be thought of as playing an analogous role to probability densities in classical statistics. In this work we mainly use pure quantum states, which are rank-1 quantum density operators and are equivalent to projections along a unit vector [58, 59, 60].
@@
Every classical observable $f\in L^{\infty}(X,\mu;\mathbb{R})\subset H$ is mapped to a selfadjoint multiplication operator $A_{f}\in B(H)$, $A_{f}g=fg$ for all $g\in H$ referred to as a quantum observable. Additionally, every classical probability density $p\in L^{1}(X,\mu;\mathbb{R})$, $\sqrt{p}\in H$ is mapped to the quantum density operator $\rho_{p}\in Q(H)$, $\rho_{p}=\langle\sqrt{p},\cdot\rangle\sqrt{p}$ representing a pure quantum state. The expected value of observable $A\in B(H)$ when the quantum system is in state $\rho\in Q(H)$ is given by $\mathbb{E}_{\rho}A=\operatorname{tr}(\rho A)\in\mathbb{R}$ which is connected to the theory of quantum measurement [58, 59, 60].
@@
We define the surrogate unresolved state space $\tilde{\mathcal{Y}}=Q(H)$ and use a quantum state $\rho\in Q(H)$ to model the state of the unresolved degrees of freedom of the dynamics. In addition, we define the operators $A_{j}\in B(H)$, $j\in\{0,\ldots,d-1\}$ to model $d$ observables of the system’s unresolved state, which are constructed by mapping the true flux functions $z_{j}\in H$ to multiplication operators. With these definitions in place we introduce the surrogate flux functions $\tilde{z}_{j}\colon Q(H)\rightarrow\mathbb{R}$, $j\in\{0,\ldots,d-1\}$ with action
$\tilde{z}_{j}(\rho)=\mathbb{E}_{\rho}(A_{j})=\operatorname{tr}(\rho A_{j}).$
@@
In other words we use the space of quantum states $\tilde{\mathcal{Y}}=Q(H)$ to model the unresolved degrees of freedom of the dynamics and the framework of quantum measurement to compute their fluxes.
@@
In addition to evaluating the required flux terms we need to be able to evolve a quantum state under the system’s dynamics and condition it based on observed dynamical behavior. In the classical case, the square root of probability densities $\sqrt{p}\in H$ evolves according to the unitary group of transfer operators $P^{t}\colon H\to H$, $P^{t}\sqrt{p}=\sqrt{p}\circ\Phi^{-t}$ with $t\geq 0$. Accordingly, the evolution of quantum states $\rho\in Q(H)$ is governed by the induced transfer operator $\mathcal{P}^{t}\colon Q(H)\to Q(H)$, $\mathcal{P}^{t}\rho=P^{t}\rho P^{-t}$. In analogy to bayesian conditioning of classical probability densities, we can condition a quantum state $\rho\in Q(H)$ using a quantum effect operator $e\in B(H)$, $0\leq e\leq\operatorname{Id}$
$\rho|_{e}=\frac{\sqrt{e}\rho\sqrt{e}}{\mathbb{E}_{\rho}e}$ (2)
@@
A quantum effect operator can be thought of as the quantum analogue of an indicator function used in classical probability to represent an event, namely a set in the event space. For instance, for a given set $F\subset X$ we can build the indicator function $1_{F}\in H$ taking the value one on $F$ and zero elsewhere. The quantum effect operator $e_{F}\in B(H)$ can then be defined as multiplication by $1_{F}$, namely $e_{F}f=1_{F}f$ for all $f\in H$.
@@ 3.2 Closure modeling
We build a finite dimensional subspace $H^{L}\subset H$ using the leading $L$ elements of an orthonormal basis $\{\phi_{\ell}\}_{\ell=0}^{\infty}$ of $H$. To obtain such a basis we define a positive kernel integral operator $K\colon H\to H$
@@
using a positive, symmetric and uniformly continuous kernel function $\kappa\in C_{b}(X\times X;\mathbb{R})\cap L^{2}(X\times X,\mu\times\mu;\mathbb{R})$ where $C_{b}$ denotes the space of bounded continuous functions. Operator $K$ is compact and selfadjoint, so its eigenfunctions form an orthonormal basis of $H$; we write $K\phi_{\ell}=\lambda_{\ell}\phi_{\ell}$ where the eigenvalues $\lambda_{\ell}$ are nonnegative and ordered in decreasing order. We define $H^{L}$ as the linear span of the leading $L$ eigenfunctions of operator $K$. Having built $H^{L}$ we project operators acting on $H$ to operators acting on $H^{L}$; that is, $A\in B(H)$ is approximated by $A_{L}=\operatorname{proj}_{L}A\operatorname{proj}_{L}$ where $\operatorname{proj}_{L}\in B(H)$ is the orthogonal projection onto $H^{L}$. This is done for all operators needed for the quantum mechanical embedding of the original classical dynamics.
@@
As explained earlier, to build our closure scheme we set $\tilde{\mathcal{Y}}=Q(H^{L})$ and use quantum states to model the unresolved degrees of freedom of the surrogate dynamics $\tilde{\Phi}^{t}$ from (1). We now describe the calculation of the fluxes needed for evolving the resolved state under $\tilde{\Phi}_{\mathcal{X}}^{t}$, and the surrogate map $\tilde{\Phi}_{\mathcal{Y}}^{t}$ used to evolve the quantum state representing the unresolved state. To do so, we use a discrete time version of the above flow maps, denoting them by $\tilde{\Phi}_{\mathcal{X}}^{n\Delta t}$ and $\tilde{\Phi}_{\mathcal{Y}}^{n\Delta t}$ with $n\in\mathbb{N}$ and timestep $\Delta t>0$.
@@
Given the resolved system state $\hat{x}_{n}\in\mathcal{X}$ at time $t_{n}=n\Delta t$ we use a quantum state $\rho_{n}\in Q(H^{L})$ to represent the unresolved state and $d$ quantum observables $A_{j}\in B(H^{L})$, $j\in\{0,\ldots,d-1\}$. Each observable $A_{j}$ is a projected multiplication operator defined using the corresponding true flux function $z_{j}\in H$. We seek to compute state $\hat{x}_{n+1}=\tilde{\Phi}_{\mathcal{X}}(\hat{x}_{n},\tilde{z}(\rho_{n}))$ using the collection of surrogate flux functions $\tilde{z}=\{\tilde{z}_{j}\}_{j=0}^{d-1}$ with $\tilde{z}_{j}\colon Q(H^{L})\to\mathbb{R}$. Each surrogate flux term $\tilde{z}_{j}(\rho_{n})$ is computed as the expectation value of observable $A_{j}$ with respect to quantum state $\rho_{n}$
$\tilde{z}_{j}(\rho_{n})=\mathbb{E}_{\rho_{n}}(A_{j})=\operatorname{tr}{(\rho_{n}A_{j})}.$
The resolved state is then updated by $\hat{x}_{n+1}=\tilde{\Phi}_{\mathcal{X}}^{\Delta t}(\hat{x}_{n},\tilde{z}(\rho_{n}))$.
@@
Next we define the surrogate unresolved dynamics map $\tilde{\Phi}_{\mathcal{Y}}^{\Delta t}$ used to update the quantum state $\rho_{n+1}=\tilde{\Phi}_{\mathcal{Y}}^{\Delta t}(\hat{x}_{n},\rho_{n})$. The quantum state update consists of two steps, a prediction and a correction step. Given the current quantum state $\rho_{n}$ we first use the projected transfer operator $\mathcal{P}^{L}\colon Q(H^{L})\rightarrow Q(H^{L})$ to compute the prior state
$\rho_{n+1}^{\prime}=\frac{\mathcal{P}^{L}\rho_{n}}{\operatorname{tr}{(\mathcal{P}^{L}\rho_{n})}}.$
@@
Renormalizing the quantum state after acting with $\mathcal{P}^{L}$ is required because the projected operator $\mathcal{P}^{L}$ is generally not unitary. The mapping $\rho_{n}\mapsto\rho_{n+1}^{\prime}$ corresponds to the prediction step. The correction step involves using the updated resolved state $\hat{x}_{n+1}$ to update our prediction $\rho_{n+1}^{\prime}$. To do so, we use state $\hat{x}_{n+1}$ to build a quantum effect operator $e_{n+1}\in B(H^{L})$; we refer to the mapping $\hat{x}_{n+1}\mapsto e_{n+1}$ as the feature map. Using the quantum Bayes rule (2) we condition the prior state $\rho_{n+1}^{\prime}$ to compute the posterior state
$\rho_{n+1}=\rho_{n+1}^{\prime}|_{e_{n+1}}=\frac{\sqrt{e_{n+1}}\rho_{n+1}^{\prime}\sqrt{e_{n+1}}}{\operatorname{tr}{(\sqrt{e_{n+1}}\rho_{n+1}^{\prime}\sqrt{e_{n+1}}})}.$
@@
The above two steps make up the surrogate unresolved dynamics $\tilde{\Phi}_{\mathcal{Y}}^{\Delta t}$. Although not unitary, operator $\mathcal{P}^{L}$ is trace nonincreasing which means that it can be interpreted as governing the evolution of an open quantum system $H^{L}\subset H$. Under this light, the reduction of $\operatorname{tr}(\rho)$ when acting with $\mathcal{P}^{L}$ can be viewed as decoherence or leakage of density due to interaction with the environment [58, 59, 60].
@@
In the diagram $\operatorname{Id}$ denotes the identity map.
@@ 4 Closure of partial differential equations
In Section 4.1 we define the appropriate Hilbert space $H$ and describe the process of embedding the classical spatiotemporal dynamics into a quantum mechanical system in $H$. In Section 4.2 we introduce the kernel used to discretize (compress) the embedded dynamics to a finite dimensional subsystem. Section 4.3 explores the symmetry invariance property of the derived eigenfunction basis. Finally, Sections 4.4 and 4.5 present the data driven formulation of the extended QMCl scheme that can be realized numerically.
@@ 4.1 Closure modeling
As explained in Section 2 we consider spatiotemporal dynamics $\Phi^{t}\colon X\rightarrow X$ possessing an invariant probability measure $\mu$ and a generally infinite dimensional state space $X\subset L^{2}(S,\nu;\mathbb{R}^{d})$ where $S$ denotes a bounded spatial domain. We assume the state space decomposition $X=\mathcal{X}\times\mathcal{Y}$ where $\mathcal{X}$ is the finite dimensional resolved state space and $\mathcal{Y}$ the unresolved state space.
To provide a general definition of the resolved state space $\mathcal{X}\subset X$ we consider the measure space $(S,\mathcal{A},\nu)$ where $\mathcal{A}$ denotes the Borel $\sigma$-algebra on $S$. We form the subalgebra $\mathcal{A}_{M}\subset\mathcal{A}$ generated by a finite collection of $M\in\mathbb{N}$ disjoint subsets $S_{m}\subset S$, $m\in\{0,\ldots,M-1\}$ such that $\cup_{m=0}^{M-1}S_{m}=S$. Next we define measure $\nu_{M}$ as the restriction of $\nu$ to $\sigma$-algebra $\mathcal{A}_{M}$. The resolved state space is defined as the $dM$-dimensional subspace $\mathcal{X}=L^{2}(S,\nu_{M};\mathbb{R}^{d})\cap X$. As a Hilbert space, $\mathcal{X}$ is isomorphic to $\mathbb{R}^{dM}$ with a scaled euclidean inner product. Each subset $S_{m}\subset S$ used to generate subalgebra $\mathcal{A}_{M}$ can also be represented by its center point $s_{m}\in S_{m}$. We denote by $S_{M}=\{s_{m}\}_{m=0}^{M-1}\subset S$ the collection of those center points. Every equivalence class of functions $\tilde{h}\in L^{2}(S,\nu_{M};\mathbb{R}^{d})$ has a piecewise constant representative $h=\sum_{j=0}^{M-1}\alpha_{j}1_{S_{j}}$ where $1_{S_{j}}$ denotes the indicator function of set $S_{j}\in\{S_{m}\}_{m=0}^{M-1}$ and $\alpha_{j}\in\mathbb{R}^{d}$ the corresponding coefficient.
As our quantum mechanical Hilbert space $H$ we consider the space of vector valued observables $H=L^{2}(X,\mu;H_{S})$ with $H_{S}=L^{2}(S,\nu_{M};\mathbb{C})$. Introducing the product state space $\Omega=X\times S$ and measure $\sigma=\mu\times\nu_{M}$ we define $H_{\Omega}=L^{2}(\Omega,\sigma;\mathbb{C})$ and note the following Hilbert space isomorphisms
@@
where $H_{X}=L^{2}(X,\mu;\mathbb{C})$. Namely, for every vector valued observable $\vec{f}\in H$ there is a corresponding complex valued $f\in H_{\Omega}$ such that for almost every $x\in X$ and $s\in S$ $f(x,\cdot)=\vec{f}(x)\in H_{S}$ and $f(x,s)=\vec{f}(x)(s)\in\mathbb{C}$. The motivation for our choice of space $H$ is our interest in observables that can effectively capture spatiotemporal dynamics; that is, patterns of evolution that have both a temporal dependence $x\in X$ and a spatial dependence $s\in S$. Given an observable $\vec{f}\in H$ the map $t\mapsto\vec{f}(\Phi^{t}(x))$ represents a spatiotemporal pattern generated by the dynamics $\Phi^{t}$ [65].
The definitions of quantum observables, quantum effects and the induced transfer operator carry over from Section 3 after using the new definition of Hilbert space $H$. To model the unresolved degrees of freedom we introduce the surrogate unresolved state space $\tilde{\mathcal{Y}}=L^{2}(S,\nu_{M};B(H))$. Namely, we use a field of quantum states over $S$, $\rho\in\tilde{\mathcal{Y}}$ with $\operatorname{ran}{\rho}\subseteq Q(H)\subset B(H)$, where each quantum state $\rho(s)$ represents the unresolved degrees of freedom of the dynamics at point $s\in S$ of the spatial domain. This allows us to compute fluxes that vary over the spatial domain $S$ for each quantum observable $A_{j}\in B(H)$, $j\in\{0,\ldots,d-1\}$. More specifically, we define the corresponding $d$ surrogate flux functions $\tilde{z}_{j}\colon\tilde{\mathcal{Y}}\to H_{S}$ with action
$\tilde{z}_{j}(\rho)(s)=\operatorname{tr}(\rho(s)A_{j})\in\mathbb{R}$
yielding $d$ flux terms at each spatial point.